Can taking chess lessons and playing chess improve memory? The online article described a study in...

  1. Can taking chess lessons and playing chess improve memory? The online article described a study in which sixth-grade students who had not previously played chess participated in a program in which they took chess lessons and played chess daily for 9 months. Each student took a memory test (the Test of Cognitive Skills) before starting the chess program and again at the end of the 9-month period. The table below shows the results for 12 randomly selected students.

Student

1

2

3

4

5

6

7

8

9

10

11

12

Pre-chess Memory Score

410

610

640

675

500

780

610

625

450

720

675

670

Post-chess Memory Score

550

670

850

650

710

795

700

795

690

710

785

675

Memory Score Difference

Test the claim at the α = 0.05 level of significance that students who participated in the chess program achieve higher memory scores after completion of the program.

  1. Fill in the last row of the table with the differences in memory.
Answer #1

Solution:-

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: ud = 0

Alternative hypothesis: ud < 0

Note that these hypotheses constitute a one-tailed test.

Formulate an analysis plan. For this analysis, the significance level is 0.05.

Analyze sample data. Using sample data, we compute the standard deviation of the differences (s), the standard error (SE) of the mean difference, the degrees of freedom (DF), and the t statistic test statistic (t).

Student 1 2 3 4 5 6 7 8 9 10 11 Pre-chess Post-chess d=B-A (d-dbar)^2 410 550 - 140 1501.5625 610 670 -60 1701.5625 640 850 -

s = \sqrt{\frac{\sum (d-\bar{d})^{2}}{n-1}}

s = 93.4472

SE = s / sqrt(n)

S.E = 26.9759

DF = n - 1 = 12 -1

D.F = 11

t = [ (x1 - x2) - D ] / SE

t = - 3.75

where di is the observed difference for pair i, d is mean difference between sample pairs, D is the hypothesized mean difference between population pairs, and n is the number of pairs.

Since we have a one-tailed test, the P-value is the probability that a t statistic having 11 degrees of freedom is less than - 3.75.

P-value = P(t < - 3.75)

Use the t-value calculator for finding p-values.

P-value = 0.002

Interpret results. Since the P-value (0.002) is less than the significance level (0.05), we have to reject the null hypothesis.

Reject H0. From the above test we have sufficient evidence in the favor of the claim that students who participated in the chess program achieve higher memory scores after completion of the program.